OpenStudy (anonymous):

Find three ordered pairs that satisfy the equation. y = 5 x - 8

OpenStudy (anonymous):

:o im not good in this ask @ganeshie8 or @nphuongsun93

OpenStudy (anonymous):

o put x = any three numbers and find y. for example x = 1, y = -3. this is a pair

OpenStudy (lgbasallote):

the most common thing to do is let x = 0, 1 and -1 then solve for y in each case when x = 0 y = 5(0) - 8 y = 0 - 8 y = -8 so when x = 0, y = 8. therefore the ordered pair is (0,-8) when x = -1 y = 5(-1) - 8 y = -5 - 8 y = -13 so when x = -1, y = -13 therefore the ordered pair would be (-1, -13) when x = 1 y = 5(1) - 8 y = 5 - 8 y = -3 so when x = 1, y = -3. therefore the ordered pair would be (1,-3)